OpenStudy (mayankdevnani):

How can we find out the value of sin18° ??? (without using calculator)

OpenStudy (mayankdevnani):

anyone???? please don't shy, answer it....

OpenStudy (dumbcow):

i know the exact answer but not sure how to come up with it....i dont see any combination of half angles , sum/difference angles to get 18 it may be done with a minimal polynomial maybe

OpenStudy (mayankdevnani):

so can you help me???

OpenStudy (anonymous):

we need to use0°, 30°, 45°, 90° etc example, 18°=x. we can use 90°, right? 90°= 5.18° mean 5x. sin 5x= sin 90° sin(x+4x)=1 continue it.

OpenStudy (dumbcow):

http://www.maths.surrey.ac.uk/hosted-sites/R.Knott/Fibonacci/simpleTrig.html its the golden ratio!!

OpenStudy (mayankdevnani):

i can't understand. @Machida

OpenStudy (anonymous):

letx=18* 5x=90* 3x=90*-2x cos3x=cos(90*-2x)=sin2x 4cos^3x-3cosx=2sinxcosx since cosx is not=0 4cos^2x-3=2sinx 4(1-sin^2x)-3=2sinx 4sin^2x+2sinx-1=0 sinx=[-2+/-rt(2^2-4*4(-1))]/2*4 sin18*=(-1+/-rt5)/4 18* is an acuteangle.so sin18*>0 therefore sinx=(-1+rt5)/4 hope u understand !!!!